OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? (Points : 1) 18 19 20 21

Directrix (directrix):

19

OpenStudy (anonymous):

let #1=x; #2=y x+y=51 x-y=13 solve for x and y

OpenStudy (anonymous):

38?

OpenStudy (anonymous):

x=y+13 y+y+13=51 2y=38 y=19 x=19+13 x=32 smaller = 19